You conduct 73 Bernoulli trails with probability 0.718 of success. What is the probability that you obtain exactly 37 successes? Round your answer to three decimal places, even if the third decimal place is 0
Answer #1

The following information is provided:

The population proportion of success is p=0.718, also, 1−p=1−0.718=0.282, and the sample size is n=73. We need to compute Pr(X=37)

This implies that

which means that the probability we are looking for is

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!